# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 1.1. Setting of the problem
In this setting, the typical behavior of the spectrum of $X_{N}=G_{N}+B_{N}$ is well known. The empirical measure of the eigenvalues of $X_{N}$ converges towards the free convolution $\sigma_{t}\boxplus\nu$ where $\sigma_{t}$ is the semicircular distribution with variance $t$[VDN92, BV93]. This probability measure was described in detail in [Bia97]. If we denote by $G_{\mu}$ the Stieltjes transform of a probability measure $\mu$ on $\mathbb{R}$ given, for $z$ outside the support of $\mu$, by
$G_{\mu}(z)=\int\frac{\mathrm{d}\mu(x)}{z-x}\,,$
with
$\Omega_{\nu,t}:=\{u+iv\in\mathbb{C}^{+}:v>v_{t}(u)\}$ (1.1)
where for all $u\in\mathbb{R}$,
$G_{\nu\boxplus\sigma_{t}}(z)=G_{\nu}(\omega_{\nu,t}(z))\,.$
When $-\Lambda$ is large enough, it is well-known that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}=G_{N}+B_{N}$ escapes the bulk, becoming smaller than $\ell_{\nu,t}$, a phenomenon called the Ben-Arous-Baik-Peche (BBP) transition. Throughout the paper, we denote
and
The following convergence result was proved in [CDMFF11]:
Theorem 1.1.
Our aim is to study the large deviations for the law of $\lambda_{1}(X_{N})$. Such a question was addressed by Ben McKenna [McK21] in the case where $B_{N}$ has no outlier, namely when $\Lambda=\ell_{\nu}$. The strategy used in [McK21] is based on a tilt by spherical integrals introduced in [GH20] and subsequently developed in diverse contexts in [GH22, CDG23, DHG24, GM20, AGH21]. However, this strategy can only be used to study large deviations below the limiting outlier $\rho_{\nu,t}^{\Lambda}$. On the other hand, [Mai07] studied the case where $\nu=\delta_{0}$, namely when $B_{N}$ has only one non trivial eigenvalue $\Lambda$. Large deviations were then obtained thanks to an exact estimation of the density due to the asymptotics of spherical integrals derived in [GM05]. In this paper, we prove a large deviation principle for $\lambda_{1}(X_{N})$, unifying the results of [Mai07] (where $\nu=0$) and [McK21] (where no outlier is permitted, namely $\Lambda=\ell_{\nu}$). Our approach is new and consists of estimating the density of the law of the smallest eigenvalue by exhibiting a fixed point functional equation for the large deviation rate function.
@@
Our study is motivated by a problem arising in the theory of spin glasses. It is well known in the physics literature that the equilibrium properties of the $p$-spin glass model are related to the topology of the energy landscape of a random function called the Thouless-Anderson-Palmer (TAP) free energy [TAP77]. In the case of the pure Ising $p$-spin glass, the Hessian of the TAP free energy contains a projector term, which may produce a negative eigenvalue at low enough temperature [Ple99, BM80, PR04]. Therefore, computing the (annealed) number of local minima requires studying the large deviations of a random matrix of the form $G_{N}+D_{N}-\Lambda uu^{T}$, where $G_{N}$ is a GOE matrix, $D_{N}$ is a diagonal matrix, $\Lambda\in\mathbb{R}$, and $u\in\mathbb{R}^{N}$. This corresponds to the model considered in this paper.
Several other works have studied large deviation principles (LDPs) for the extreme eigenvalues of GOE matrices. The article [Mai07], which examined a rank-one deformation of the GOE was extended in [BGGM12] to finite rank deformations, and in [BG20], which proved an LDP for the joint distribution of extreme eigenvalues and their associated eigenvectors. By leveraging the computation of spherical integrals, [GH20] and [GM20] investigated the extreme eigenvalues of sharp subgaussian matrices and free convolution models. In [McK21], an LDP was obtained for $G_{N}+B_{N}$ where $B_{N}$ is a diagonal matrix without outliers with a rate function expressed variationally.
The smallest eigenvalue of Wigner matrices is highly sensitive to the entry distribution’s tail behavior so that even the speed of its large deviation principle depends on this tail. For distributions with tails heavier than Gaussian, an LDP was established in [Aug16], showing the speed and the rate function depend only on the tail decay rate. Sharp sub-Gaussian and general sub-Gaussian distributions have been widely studied. Sharp sub-Gaussian distributions have tails asymptotically equivalent to Gaussian, resulting in a rate function matching the one of the GOE [BC14, AGH21]. In contrast, general sub-Gaussian distributions may exhibit a transition to a non-universal rate function, indicating eigenvector localization phenomena [CDG23]. Recently, an LDP for the largest eigenvalue of random matrices with a variance profile was established, highlighting the impact of the variance profile on the rate function [Hus22, DHG24].
@@ 1.2. Main result
The GOE of size $N\times N$ of variance $t>0$ is the symmetric random matrix $(G_{ij})_{1\leq i,j\leq N}$ where for each $i\leq j$ and $k\leq l$ such that $(i,j)\neq(k,l)$, $G_{ij}$ and $G_{kl}$ are independent and such that for each $i\leq j$,
@@
Assumption (4) should not be necessary. Moreover, our approach extends readily to the case where $G_{N}$ is a GUE matrix instead of a GOE matrix.
@@ 1.3. Notation
    $\ell_{\nu,t}^{\Lambda}=\begin{cases}\rho_{\nu,t}^{\Lambda}&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$}\\ \ell_{\nu,t}&\text{if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$}.\end{cases}$
  • For every $\nu\in\mathcal{P}(\mathbb{R})$ and $\lambda\in[H_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$, we denote $\omega_{\nu,t}^{*}(\lambda)$ the unique solution in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$ of the equation $H_{\nu,t}(\gamma)=\lambda$.
@@
  • For each $p\geq 1$, we denote $\Delta^{p}$ the $p$-simplex
@@ 1.4. Heuristics
We assume in these heuristics that the matrix $B_{N}$ has eigenvalues $\Lambda:=\eta_{0}<\eta_{1}<\cdots<\eta_{p}$ with multiplicities $1$, $N_{1},\ldots,N_{p}$, such that for each $i=1,\ldots,p$, $N_{i}/N$ converges to $\alpha_{i}>0$. We outline the arguments to derive the weak large deviation principle, namely to estimate the probability that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}=G_{N}+B_{N}$ is close to a given ${x\leq\ell_{\nu,t}}$. Let $\varepsilon>0$.
The probability that the smallest eigenvalue $\lambda_{1}(X_{N})$ of ${X_{N}}$ lies in $(x-\varepsilon,x+\varepsilon)$ is given by
@@
where $\mathrm{d}X_{N}$ is the Lebesgue measure on symmetric matrices of size $N\times N$ and $Z_{N}^{t}$ is the normalizing constant. Let us more generally denote by $\lambda_{1}\leq\cdots\leq\lambda_{N}$ the eigenvalues of $X_{N}$ and ${(v_{1},\ldots,v_{N})}$ the associated unitary eigenvectors.
One can split the term in the density into
@@
where $D(\lambda)=\mbox{diag}(\lambda_{2},\ldots,\lambda_{N})$ and $O$ follows the Haar measure on $v_{1}^{\perp}$. For every $\lambda\in\{\eta_{0},\ldots,\eta_{p}\}$, let us denote $E_{\lambda}$ the associated eigenspace of $B_{N}$. We observe that $\langle v_{1},B_{N}v_{1}\rangle$ and $\langle v_{1},B_{N}^{2}v_{1}\rangle$ only depend on ${Y_{N}:=\{\|v_{1}\|_{E_{\eta_{k}}}^{2},k=0,1,\ldots,p\}}$ so that the first term in the above density is a polynomial $L_{\nu,t}^{\Lambda}$ in $Y_{N}$ (defined in (2.16)).
Fix $Y=(y_{0},\ldots,y_{p})\in\Delta^{p+1}$, where $\Delta^{p+1}$ is the $(p+1)$-simplex (1.14). Restricting the domain of integration to vectors $v_{1}\in\mathbb{R}^{N}$ such that $Y_{N}$ is close to $Y$ and using the large deviation principle for $Y_{N}$ when $v_{1}$ follows the uniform law on the sphere, we roughly obtain, since the empirical measure of the eigenvalues of $B_{N}$ is close to $\nu:=\sum_{i=1}^{p}\alpha_{i}\delta_{\eta_{i}}$, the lower bound
$\liminf_{N\to\infty}\frac{1}{N}\log\frac{1}{Z_{N}^{t}}\int_{\lambda_{1}(X_{N})\in(x-\varepsilon,x+\varepsilon)}e^{-\frac{N}{4t}\operatorname{Tr}((X_{N}-B_{N})^{2})}\,\mathrm{d}X_{N}\gtrsim-J_{\nu,t}^{\Lambda}(x,Y)-\int\log|\lambda-x|\,\mathrm{d}(\nu\boxplus\sigma_{t})(\lambda)\\ +\frac{x^{2}}{4t}+C_{t}+\liminf_{N\to\infty}\frac{1}{N}\log\frac{1}{Z_{N-1}^{t(N-1)/N}}\int_{\lambda_{2}\geq x}e^{-\frac{N}{4t}\operatorname{Tr}((p_{v_{1}}^{\perp}(X_{N}-B_{N})p_{v_{1}}^{\perp})^{2})}\,\mathrm{d}X_{N},$ (1.15)
where $C_{t}$ is a finite constant and
@@
By Courant-Fischer’s theorem, when ${Y_{N}\simeq Y}$, the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has a unique outlier approximately equal to $\Lambda_{1}:=\Phi(\Lambda,Y)$, which is defined as the smallest solution of the equation
@@
Moreover, the spectrum of $B_{N}^{\prime}=p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ converges to ${\nu=\sum_{k=1}^{p}\alpha_{k}\delta_{\eta_{k}}}$. As a consequence, we see that the last term in (1.15) is also the probability of deviation of the smallest eigenvalue $\lambda_{2}$ of a $(N-1)\times(N-1)$ matrix given by the sum of a GOE matrix and a matrix $B_{N}^{\prime}$ with empirical measure of the eigenvalues converging to $\nu$ and with outlier $\Phi(\Lambda,Y)$. We set
@@
As a consequence of the previous considerations, we obtain a functional inequality
$-F_{\nu,t}^{-}(\Lambda,x)\gtrsim\sup_{Y\in\Delta^{p+1}}\Bigl{\{}J_{\nu,t}^{\Lambda}(x,Y)-C_{t}-\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)+\frac{x^{2}}{4t}-\inf_{y\geq x}F_{\nu,t}^{-}(\Phi(\Lambda,Y),y)\Bigr{\}},$
@@
$-F_{\nu,t}^{+}(\Lambda,x)\lesssim\sup_{Y\in\Delta^{p+1}}\Bigl{\{}J_{\nu,t}^{\Lambda}(x,Y)-C_{t}-\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)+\frac{x^{2}}{4t}-\inf_{y\geq x}F_{\nu,t}^{+}(\Phi(\Lambda,Y),y)\Bigr{\}}\,.$
@@
$I_{\nu,t}^{\Lambda}(\lambda)=\inf_{Y\in\Delta^{p+1}}\Bigr{\{}J_{\nu,t}^{\Lambda}(\lambda,Y)-C_{t}-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}\,.$

### v2

@@ 1.1. Setting of the problem
In this setting, the typical behavior of the spectrum of $X_{N}=G_{N}+B_{N}$ is well known. The empirical measure of the eigenvalues of $X_{N}$ converges towards the free convolution $\sigma_{t}\boxplus\nu$ where $\sigma_{t}$ is the semicircular distribution with variance $t$[VDN92, BV93]. This probability measure was described in detail in [Bia97]. Moreover, it was proved in [CDMFF11] that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}$ converges towards a limit $\ell^{\Lambda}_{\nu,t}$ which is equal to the left boundary $\ell_{\nu,t}$ of the support of $\sigma_{t}\boxplus\nu$ if and only if $\Lambda$ is above some threshold, a phenomenon called the Ben-Arous-Baik-Péché (BBP) transition (see Theorem 1.1).
@@
Our aim is to study the large deviations for the law of $\lambda_{1}(X_{N})$, the smallest eigenvalue of $X_{N}$. Our study is motivated by a problem arising in the theory of spin glasses. It is well known in the physics literature that the equilibrium properties of the $p$-spin glass model are related to the topology of the energy landscape of a random function called the Thouless-Anderson-Palmer (TAP) free energy [TAP77]. In the case of the pure Ising $p$-spin glass, the Hessian of the TAP free energy contains a projector term, which may produce a negative eigenvalue at low enough temperature [Ple99, BM80, PR04]. By the Kac-Rice formula, computing the (annealed) number of local minima therefore requires studying the large deviations of the smallest eigenvalue of a random matrix of the form considered in this paper.
Several other works have studied large deviation principles (LDPs) for the extreme eigenvalues of GOE matrices. The article [Mai07], which examined a rank-one deformation of the GOE, was extended in [BGGM12] to finite rank deformations, and in [BG20], which proved an LDP for the joint distribution of extreme eigenvalues and their associated eigenvectors. The article [GH20] investigated the extreme eigenvalues of Wigner matrices, namely self-adjoint random matrices with entries independent above the diagonal. They showed that if the entries are sharp subgaussian in the sense that their Laplace transform is bounded above by the Laplace transform of the Gaussian variable with the same variance, the large deviations are the same as for Gaussian entries. This came as a surprise, as large deviations for the smallest eigenvalue of Wigner matrices are generally very sensitive to the distribution of the entries, so that even their speed depends on their tail. For distributions with tails heavier than Gaussian, an LDP was established in [Aug16], showing the speed and the rate function depends on the tail decay rate. In contrast, the large deviations when the entries follow general sub-Gaussian distributions have the same speed as for Gaussian variables, but their rate function is non-universal and in general is different from the Gaussian one below a certain threshold, a transition indicating eigenvector localization phenomena [AGH21, CDG23]. Large deviations were also investigated for more structured matrices. They were first studied for the sum of two unitarily invariant models [GM20] and more recently for matrices with a variance profile, highlighting the impact of the variance profile on the rate function [Hus22, DHG24]. In [McK21], an LDP was obtained for $G_{N}+B_{N}$ where $B_{N}$ is a diagonal matrix without outliers, namely when $\Lambda=\ell_{\nu}$.
The strategy used in [McK21] is based on a tilt by spherical integrals introduced in [GH20] and subsequently developed in diverse contexts in [GH22, CDG23, DHG24, GM20, AGH21]. However, this strategy can only be used to study large deviations below the limiting outlier $\ell_{\nu,t}^{\Lambda}$. On the other hand, [Mai07] studied the case where $\nu=\delta_{0}$, namely when $B_{N}$ has only one non trivial eigenvalue $\Lambda$. Large deviations were then obtained thanks to an exact estimation of the density due to the asymptotics of spherical integrals derived in [GM05]. In this paper, we prove a large deviation principle for $\lambda_{1}(X_{N})$, unifying the results of [Mai07] (where $\nu=0$) and [McK21] (where no outlier is permitted, namely $\Lambda=\ell_{\nu}$). Our approach is new and consists of estimating the density of the law of the smallest eigenvalue by exhibiting a fixed point functional equation for the large deviation rate function.
@@ 1.2. Main result
The GOE of size $N\times N$ of variance $t>0$ is the symmetric random matrix $(G_{ij})_{1\leq i,j\leq N}$ where for every $i\leq j$ and $k\leq l$ such that $(i,j)\neq(k,l)$, $G_{ij}$ and $G_{kl}$ are independent and such that for every $i\leq j$,
@@
Under this assumption, the empirical measure and the smallest eigenvalue of $X_{N}=B_{N}+G_{N}$ converge almost surely. We now describe their limits more precisely. If we denote by $G_{\mu}$ the Stieltjes transform of a probability measure $\mu$ on $\mathbb{R}$ given, for $z$ outside the support of $\mu$, by
$G_{\mu}(z)=\int\frac{\mathrm{d}\mu(x)}{z-x}\,,$
with
$\Omega_{\nu,t}:=\{u+iv\in\mathbb{C}^{+}:v>v_{t}(u)\}$ (1.5)
where for all $u\in\mathbb{R}$,
$G_{\nu\boxplus\sigma_{t}}(z)=G_{\nu}(\omega_{\nu,t}(z))\,.$
When $-\Lambda$ is large enough, it is well-known that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}=G_{N}+B_{N}$ escapes the bulk, becoming smaller than $\ell_{\nu,t}$. Throughout the paper, we denote
and
The following convergence result was proved in [CDMFF11]:
Theorem 1.1.
@@
(a) $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$
(b) $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda\leq\rho_{\nu,t}^{\Lambda}$
(c) $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$ and $\lambda>\rho_{\nu,t}^{\Lambda}$
Figure 1. Graph of $H_{\nu,t}:x\in(-\infty,\ell_{\nu})\mapsto x+tG_{\nu}(x)$. Figures $A$ and $B$ correspond to the first case, Figure $C$ to the second case of (1.7).
@@
Then, the main result of our paper is the following large deviation principle.
@@
Assumption 1.1.(4) should not be necessary. Moreover, our proof extends readily to the case where $G_{N}$ is a GUE matrix instead of a GOE matrix.
@@ 1.3. Notation
    $\ell_{\nu,t}^{\Lambda}=\begin{cases}\rho_{\nu,t}^{\Lambda}&\text{if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$}\\ \ell_{\nu,t}&\text{if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$}.\end{cases}$ (1.14)
  • For every $\nu\in\mathcal{P}(\mathbb{R})$ and $\lambda\in[H_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$, we denote $\omega_{\nu,t}^{*}(\lambda)$ the unique solution $\gamma$ of
    $H_{\nu,t}(\gamma)=\lambda,\quad\gamma\in[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu}).$ (1.15)
@@
  • For every $p\geq 1$, we denote $\Delta^{p}$ the $p$-simplex
@@
  • $\lambda^{N}=(\lambda_{1},\ldots,\lambda_{N})$ denotes a $N$ dimensional vector in $\mathbb{R}^{N}$ and $\mathrm{d}\lambda^{N}=\mathrm{d}\lambda_{1}\cdots\mathrm{d}\lambda_{N}$.
@@ 1.4. Heuristics
We assume in these heuristics that the matrix $B_{N}$ has eigenvalues $\Lambda:=\eta_{0}<\eta_{1}<\cdots<\eta_{p}$ with multiplicities $1$, $N_{1},\ldots,N_{p}$, such that for every $i=1,\ldots,p$, $N_{i}/N$ converges to $\alpha_{i}>0$. We outline the arguments to derive the weak large deviation principle, namely to estimate the probability that the smallest eigenvalue $\lambda_{1}(X_{N})$ of $X_{N}=G_{N}+B_{N}$ is close to a given ${x\leq\ell_{\nu,t}}$.
Let $\varepsilon>0$. The probability that $\lambda_{1}(X_{N})$ lies in $(x-\varepsilon,x+\varepsilon)$ is given by
@@
where $\mathrm{d}X_{N}$ is the Lebesgue measure on symmetric matrices of size $N\times N$ and $Z_{N}^{t}$ is the normalizing constant. Let us denote by $\lambda_{1}\leq\cdots\leq\lambda_{N}$ the eigenvalues of $X_{N}$ and ${(v_{1},\ldots,v_{N})}$ the associated unitary eigenvectors. One can split the term in the density into
@@
where $D(\lambda)=\mbox{diag}(\lambda_{2},\ldots,\lambda_{N})$ and $O$ follows the Haar measure on $v_{1}^{\perp}$. For every $\lambda\in\{\eta_{0},\ldots,\eta_{p}\}$, let us denote $E_{\lambda}$ the associated eigenspace of $B_{N}$. We first observe that $\langle v_{1},B_{N}v_{1}\rangle$ and $\langle v_{1},B_{N}^{2}v_{1}\rangle$ only depend on ${Y_{N}:=\{\|v_{1}\|_{E_{\eta_{k}}}^{2},k=0,1,\ldots,p\}}$ so that the first term in the above density is a polynomial $-L_{\nu,t}^{\Lambda}$ in $\lambda_{1}$ and $Y_{N}$, see Lemma 2.2. Moreover, $OD(\lambda)O^{T}$ follows the law of the $(N-1)\times(N-1)$ GOE matrix $\sqrt{N/N-1}G_{N-1}$ (living in $v_{1}^{\perp}$) under the distribution
$\frac{1}{Z_{N-1}^{t(N-1)/N}}e^{-\frac{N}{4t}\operatorname{Tr}((OD(\lambda)O^{T})^{2})}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|1_{\lambda_{1}\leq\ldots\leq\lambda_{N}}\,\mathrm{d}\lambda_{2}\cdots\,\mathrm{d}\lambda_{N}\,\mathrm{d}O\,.$
Fix $Y=(y_{0},\ldots,y_{p})\in\Delta^{p+1}$, where $\Delta^{p+1}$ is the $(p+1)$-simplex (1.17). Restricting the domain of integration to vectors $v_{1}\in\mathbb{R}^{N}$ such that $Y_{N}$ is close to $Y$ and using the large deviation principle for $Y_{N}$ when $v_{1}$ follows the uniform law on the sphere (see Lemma 2.3), we roughly obtain, since the empirical measure of the eigenvalues of $X_{N}$ is close to $\nu\boxplus\sigma_{t}$ with overwhelming probability, the lower bound
$\liminf_{N\to\infty}\frac{1}{N}\log\frac{1}{Z_{N}^{t}}\int_{\lambda_{1}(X_{N})\in(x-\varepsilon,x+\varepsilon)}e^{-\frac{N}{4t}\operatorname{Tr}((X_{N}-B_{N})^{2})}\,\mathrm{d}X_{N}\gtrsim-J_{\nu,t}^{\Lambda}(x,Y){+}\int\log|\lambda-x|\,\mathrm{d}(\nu\boxplus\sigma_{t})(\lambda)\\ {-}\frac{x^{2}}{4t}-C_{t}+\liminf_{N\to\infty}\frac{1}{N}\log\frac{1}{Z_{N-1}^{t(N-1)/N}}\int_{\lambda_{2}\geq x}e^{-\frac{N}{4t}\operatorname{Tr}((X_{N-1}-B_{N-1}^{\prime})^{2})}\,\mathrm{d}X_{N-1},$ (1.18)
where $B_{N-1}^{\prime}$ is the $(N-1)\times(N-1)$ matrix living in $v_{1}^{\perp}$ where it equals $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$, $C_{t}$ is a finite constant and
@@
By Courant-Fischer’s theorem, when ${Y_{N}\simeq Y}$, the matrix $B_{N-1}^{\prime}$ has a unique outlier approximately equal to $\Lambda_{1}:=\Phi(\Lambda,Y)$, which is defined as the smallest solution in $[\Lambda,\eta_{1}]$ of the equation
@@
Note that it always exists when $y_{0}$ and $y_{1}$ are positive since the function of $\Lambda_{1}$ on the left hand side goes then goes from $\infty$- ∞ to +$\infty$+ ∞. We will see it is then unique. Moreover, the spectrum of $B_{N-1}^{\prime}$ converges to ${\nu=\sum_{k=1}^{p}\alpha_{k}\delta_{\eta_{k}}}$. As a consequence, we see that the last term in (1.18) is also the probability of deviation of the smallest eigenvalue $\lambda_{2}$ of a $(N-1)\times(N-1)$ matrix given by the sum of a GOE matrix and a matrix $B_{N-1}^{\prime}$ with empirical measure of the eigenvalues converging to $\nu$ and with outlier $\Phi(\Lambda,Y)$. We set
@@
As a consequence of the previous considerations, we obtain the functional inequality
$-F_{\nu,t}^{-}(\Lambda,x)\gtrsim\sup_{Y\in\Delta^{p+1}}\Bigl{\{}-J_{\nu,t}^{\Lambda}(x,Y)-C_{t}+\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)-\frac{x^{2}}{4t}-\inf_{y\geq x}F_{\nu,t}^{-}(\Phi(\Lambda,Y),y)\Bigr{\}}\,,$
for some constant $C_{t}$ depending only on $t$. In other words, we find
$F_{\nu,t}^{-}(\Lambda,x)\lesssim\inf_{Y\in\Delta^{p+1}}\Bigl{\{}J_{\nu,t}^{\Lambda}(x,Y)+C_{t}-\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)+\frac{x^{2}}{4t}+\inf_{y\geq x}F_{\nu,t}^{-}(\Phi(\Lambda,Y),y)\Bigr{\}}\,.$
@@
$F_{\nu,t}^{+}(\Lambda,x)\gtrsim\inf_{Y\in\Delta^{p+1}}\Bigl{\{}J_{\nu,t}^{\Lambda}(x,Y)+C_{t}-\int\log|\lambda-x|\,d(\nu\boxplus\sigma_{t})(\lambda)+\frac{x^{2}}{4t}+\inf_{y\geq x}F_{\nu,t}^{+}(\Phi(\Lambda,Y),y)\Bigr{\}}\,.$
@@
$I_{\nu,t}^{\Lambda}(\lambda)=\inf_{Y\in\Delta^{p+1}}\Bigr{\{}J_{\nu,t}^{\Lambda}(\lambda,Y)+C_{t}-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}\,.$